40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 840     Entry time: Fri Feb 24 21:59:44 2012
Author: Tara, Frank 
Type: DailyProgress 
Category: Vacuum 
Subject: upgrades installed - pumping started 

we've installed:

  • the new stack -see earlier post with measured TF
  • both copper radiation shields
    - one shield has a NiChrome heater wrapped around (the other one not as we first want to see if we need more/less power, but calculation says this one should be right)
    - three temp sensors (AD590), two on the shield with heater, one on the one without
    - those sensors won't be used for temp control, only for temp and gradient monitoring and/or safety shutdown (if required)
  • cable from D-SUB feedthrough to rack breakout terminal

window is back on and pumping down over the weekend. Thermal insulation for vacuum chamber is back in place, so we should be ready to go for a new measurement on Monday afternoon after installation of the beat breadboard

P1820701.JPG P1820756.JPG

ELOG V3.1.3-